# Semi-analytical covariance matrices for two-point correlation function for DESI 2024 data

### v1

@@ Appendix D Author affiliations
1Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
@@
2Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, CH-1290 Versoix, Switzerland
3IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
4Physics Department, Yale University, P.O. Box 208120, New Haven, CT 06511, USA
5Department of Physics & Astronomy, Ohio University, Athens, OH 45701, USA
@@
6Center for Cosmology and AstroParticle Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
@@
7Department of Astronomy, The Ohio State University, 4055 McPherson Laboratory, 140 W 18th Avenue, Columbus, OH 43210, USA
8Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
9Physics Dept., Boston University, 590 Commonwealth Avenue, Boston, MA 02215, USA
10University of Michigan, Ann Arbor, MI 48109, USA
@@
11Leinweber Center for Theoretical Physics, University of Michigan, 450 Church Street, Ann Arbor, Michigan 48109-1040, USA
@@
12Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
@@
13Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
14Instituto de Física, Universidad Nacional Autónoma de México, Cd. de México C.P. 04510, México
@@
15Department of Astronomy, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
@@
16Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Menlo Park, CA 94305, USA
17SLAC National Accelerator Laboratory, Menlo Park, CA 94305, USA
18University of California, Berkeley, 110 Sproul Hall #5800 Berkeley, CA 94720, USA
@@
19Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra Barcelona, Spain
20Departamento de Física, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio Ip, CP 111711, Bogotá, Colombia
@@
21Observatorio Astronómico, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio H, CP 111711 Bogotá, Colombia
22Department of Physics, The University of Texas at Dallas, Richardson, TX 75080, USA
@@
23Departament de Física Quàntica i Astrofísica, Universitat de Barcelona, Martí i Franquès 1, E08028 Barcelona, Spain
24Institut d’Estudis Espacials de Catalunya (IEEC), 08034 Barcelona, Spain
@@
25Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain.
@@
26Consejo Nacional de Ciencia y Tecnología, Av. Insurgentes Sur 1582. Colonia Crédito Constructor, Del. Benito Juárez C.P. 03940, México D.F. México
27Departamento de Física, Universidad de Guanajuato - DCI, C.P. 37150, Leon, Guanajuato, México
28Fermi National Accelerator Laboratory, PO Box 500, Batavia, IL 60510, USA
@@
29Department of Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
30School of Mathematics and Physics, University of Queensland, 4072, Australia
31NSF NOIRLab, 950 N. Cherry Ave., Tucson, AZ 85719, USA
@@
32Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), FR-75005 Paris, France
@@
33Departament de Física, Serra Húnter, Universitat Autònoma de Barcelona, 08193 Bellaterra (Barcelona), Spain
@@
34Laboratoire de Physique Subatomique et de Cosmologie, 53 Avenue des Martyrs, 38000 Grenoble, France
@@
35Institució Catalana de Recerca i Estudis Avançats, Passeig de Lluís Companys, 23, 08010 Barcelona, Spain
36Department of Physics and Astronomy, University of Sussex, Brighton BN1 9QH, U.K
37Department of Physics & Astronomy, University of Wyoming, 1000 E. University, Dept. 3905, Laramie, WY 82071, USA
@@
38National Astronomical Observatories, Chinese Academy of Sciences, A20 Datun Rd., Chaoyang District, Beijing, 100012, P.R. China
@@
39Instituto Avanzado de Cosmología A. C., San Marcos 11 - Atenas 202. Magdalena Contreras, 10720. Ciudad de México, México
40Department of Physics and Astronomy, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
41Waterloo Centre for Astrophysics, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
@@
42Perimeter Institute for Theoretical Physics, 31 Caroline St. North, Waterloo, ON N2L 2Y5, Canada
43Space Sciences Laboratory, University of California, Berkeley, 7 Gauss Way, Berkeley, CA 94720, USA
44Max Planck Institute for Extraterrestrial Physics, Gießenbachstraße 1, 85748 Garching, Germany
@@
45Department of Physics, Kansas State University, 116 Cardwell Hall, Manhattan, KS 66506, USA
46Department of Physics and Astronomy, Sejong University, Seoul, 143-747, Korea
@@
47Centre for Astrophysics & Supercomputing, Swinburne University of Technology, P.O. Box 218, Hawthorn, VIC 3122, Australia
48CIEMAT, Avenida Complutense 40, E-28040 Madrid, Spain
49Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
@@
50Department of Astronomy, Tsinghua University, 30 Shuangqing Road, Haidian District, Beijing, China, 100190

### v2

@@ Appendix D Author affiliations
Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
@@
Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédérale de Lausanne (EPFL), Observatoire de Sauverny, CH-1290 Versoix, Switzerland
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
Physics Department, Yale University, P.O. Box 208120, New Haven, CT 06511, USA
Department of Physics & Astronomy, Ohio University, Athens, OH 45701, USA
@@
Center for Cosmology and AstroParticle Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
@@
Department of Astronomy, The Ohio State University, 4055 McPherson Laboratory, 140 W 18th Avenue, Columbus, OH 43210, USA
Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
Physics Dept., Boston University, 590 Commonwealth Avenue, Boston, MA 02215, USA
University of Michigan, Ann Arbor, MI 48109, USA
@@
Leinweber Center for Theoretical Physics, University of Michigan, 450 Church Street, Ann Arbor, Michigan 48109-1040, USA
@@
Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
@@
Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK
Instituto de Física, Universidad Nacional Autónoma de México, Cd. de México C.P. 04510, México
@@
Department of Astronomy, School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai 200240, China
@@
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Menlo Park, CA 94305, USA
SLAC National Accelerator Laboratory, Menlo Park, CA 94305, USA
University of California, Berkeley, 110 Sproul Hall #5800 Berkeley, CA 94720, USA
@@
Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra Barcelona, Spain
Departamento de Física, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio Ip, CP 111711, Bogotá, Colombia
@@
Observatorio Astronómico, Universidad de los Andes, Cra. 1 No. 18A-10, Edificio H, CP 111711 Bogotá, Colombia
Department of Physics, The University of Texas at Dallas, Richardson, TX 75080, USA
@@
Departament de Física Quàntica i Astrofísica, Universitat de Barcelona, Martí i Franquès 1, E08028 Barcelona, Spain
Institut d’Estudis Espacials de Catalunya (IEEC), 08034 Barcelona, Spain
@@
Institut de Ciències del Cosmos (ICCUB), Universitat de Barcelona (UB), c. Martí i Franquès, 1, 08028 Barcelona, Spain.
@@
Consejo Nacional de Ciencia y Tecnología, Av. Insurgentes Sur 1582. Colonia Crédito Constructor, Del. Benito Juárez C.P. 03940, México D.F. México
Departamento de Física, Universidad de Guanajuato - DCI, C.P. 37150, Leon, Guanajuato, México
Fermi National Accelerator Laboratory, PO Box 500, Batavia, IL 60510, USA
@@
Department of Physics, The Ohio State University, 191 West Woodruff Avenue, Columbus, OH 43210, USA
School of Mathematics and Physics, University of Queensland, 4072, Australia
NSF NOIRLab, 950 N. Cherry Ave., Tucson, AZ 85719, USA
@@
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), FR-75005 Paris, France
@@
Departament de Física, Serra Húnter, Universitat Autònoma de Barcelona, 08193 Bellaterra (Barcelona), Spain
@@
Laboratoire de Physique Subatomique et de Cosmologie, 53 Avenue des Martyrs, 38000 Grenoble, France
@@
Institució Catalana de Recerca i Estudis Avançats, Passeig de Lluís Companys, 23, 08010 Barcelona, Spain
Department of Physics and Astronomy, University of Sussex, Brighton BN1 9QH, U.K
Department of Physics & Astronomy, University of Wyoming, 1000 E. University, Dept. 3905, Laramie, WY 82071, USA
@@
National Astronomical Observatories, Chinese Academy of Sciences, A20 Datun Rd., Chaoyang District, Beijing, 100012, P.R. China
@@
Instituto Avanzado de Cosmología A. C., San Marcos 11 - Atenas 202. Magdalena Contreras, 10720. Ciudad de México, México
Department of Physics and Astronomy, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
Waterloo Centre for Astrophysics, University of Waterloo, 200 University Ave W, Waterloo, ON N2L 3G1, Canada
@@
Perimeter Institute for Theoretical Physics, 31 Caroline St. North, Waterloo, ON N2L 2Y5, Canada
Space Sciences Laboratory, University of California, Berkeley, 7 Gauss Way, Berkeley, CA 94720, USA
Max Planck Institute for Extraterrestrial Physics, Gießenbachstraße 1, 85748 Garching, Germany
@@
Department of Physics, Kansas State University, 116 Cardwell Hall, Manhattan, KS 66506, USA
Department of Physics and Astronomy, Sejong University, Seoul, 143-747, Korea
@@
Centre for Astrophysics & Supercomputing, Swinburne University of Technology, P.O. Box 218, Hawthorn, VIC 3122, Australia
CIEMAT, Avenida Complutense 40, E-28040 Madrid, Spain
Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
@@
Department of Astronomy, Tsinghua University, 30 Shuangqing Road, Haidian District, Beijing, China, 100190